In the current year, Jill, age 35, received a job offer with two alternative compensation packages to choose from. The first package offers her $93,600 annual salary with no qualified fringe benefits and requires her to pay $5,300 a year for parking and to purchase life insurance at a cost of $2,800. The second package offers $81,800 annual salary, employer-provided health insurance, annual free parking (worth $500 per month), $200,000 of life insurance (purchasing on her own would have been $2,800 annually), and free flight benefits (she estimates that it will save her $6,800 per year). If Jill chooses the first package, she will purchase the health and life insurance benefits herself at a cost of $2,800 annually after taxes and spend another $6,800 in flights while traveling. Assume her marginal tax rate is 32 percent. (Use Exhibit 12-8.)

Required:

  1. a1. Which compensation package should she choose?
  2. a2. How much would she benefit in after-tax dollars by choosing this compensation package instead of the alternative package?
  3. b1. Assume the first package offers $119,000 salary with no qualified benefits instead of $93,600 salary and the other benefits and costs are the same. Which compensation package should she choose?
  4. b2. How much would she benefit in after-tax dollars by choosing this package?
Answer #1

A-1)

Value of First package = $ 93,600 (No fringe benefits)

Value of Second package = $ 81,800

Add: ) Value of parking benefits at first job (Opportunity cost) = $ 5300

Add: ) Value of flight benefits = $ 6800

Add: ) Value of insurance premium = $ 2800

Total value of Second package = $ 96,700

Second Package Net Benefit = (96,700 – 93,600) = $ 3100

Therefore Jill should choose the second package, as inclusive of the fringe benefits being provided with the salary package it is more beneficial compared to the first package.

A-2)

Value of First package = $ 93,600

Less: ) Insurance premium paid = $2800

Net Taxable Income (First package) = $ 90800

Net Tax Payable @ 32% = (90800 x 0.32) = $ 29056

Net value of First package (after tax) = (93,600 – 29056) = $ 64,544

Value of Second package = $ 81,800

Add: ) Value of parking benefits at first job (500 x 12) = $ 6000

Add: ) Value of flight benefits = $ 6,800

Add: ) Value of insurance premium = $ 2800

Net Taxable Income (Second package) = $ 97,400

Net Tax Payable @ 32% =   (97400x 0.32) = $ 31,168

Net value of Second package (after tax)   =   (97400– 31168)   =   $ 66,232

Second Package Net Benefit (after tax) = (66232 – 64,544) = $ 1688

Although due to lesser tax deductions the Net Benefit (after tax) has reduced but Jill should still choose the second package, as inclusive of the fringe benefits being provided with the salary package it is more beneficial compared to the first package.

B1)

Value of First package = $ 119,000 (No fringe benefits)

Value of Second package = $ 81,800

Add: ) Value of parking benefits at first job (Opportunity cost) = $ 5300

Add: ) Value of flight benefits = $ 6800

Add: ) Value of insurance premium = $ 2800

Total value of Second package = $ 96,700

First Package Net Benefit = (119,000 - 96,700) = $ 22,300

Therefore Jill should choose the first package, as the amount paid is higher than the salary inclusive of the fringe benefits being provided in the second package, thus it is more beneficial compared to the second package.

B2)

Value of First package = $ 119,000

Less: ) Insurance premium paid = $2800

Net Taxable Income (First package) = $ 116200

Net Tax Payable @ 32% = (116200 x 0.32) = $ 37184

Net value of First package (after tax) = (119000 – 37184) = $ 81816

Value of Second package = $ 81,800

Add: ) Value of parking benefits at first job (500 x 12) = $ 6000

Add: ) Value of flight benefits = $ 6800

Add: ) Value of insurance premium = $ 2800

Net Taxable Income (Second package) = $ 97400

Net Tax Payable @ 32% = (97400 x 0.32) = $ 31168

Net value of Second package (after tax)   =   (97400 – 31168)   =   $ 66,232

First Package Net Benefit (after tax) = (81816 - 66,232) = $ 15584

Due the higher salary and the deduction for insurance expenses the Net Benefit (after tax) for the first package is higher compared to the second package, thus Jill should choose the first package.